OpenStudy (anonymous):

Please help and show all work,step by step Write the equation of the line which passes through (2, 1) and is perpendicular to x = –2.

OpenStudy (anonymous):

1/2 = perp slope. y-1/x-2 = 1/2 2y -2 = x -2 y = 1/2x